# Tiny Models are the Computational Saver for Large Models

### v1

@@ Abstract
This paper introduces TinySaver, an early-exit-like dynamic model compression approach which employs tiny models to substitute large models adaptively. Distinct from traditional compression techniques, dynamic methods like TinySaver can leverage the difficulty differences to allow certain inputs to complete their inference processes early, thereby conserving computational resources. Most existing early exit designs are implemented by attaching additional network branches to the model’s backbone. Our study, however, reveals that completely independent tiny models can replace a substantial portion of the larger models’ job with minimal impact on performance. Employing them as the first exit can remarkably enhance computational efficiency. By searching and employing the most appropriate tiny model as the computational saver for a given large model, the proposed approaches work as a novel and generic method to model compression. This finding will help the research community in exploring new compression methods to address the escalating computational demands posed by rapidly evolving AI models. Our evaluation of this approach in ImageNet-1k classification demonstrates its potential to reduce the number of compute operations by up to 90%, with only negligible losses in performance, across various modern vision models. The code of this work will be available.
@@ 2.2.1 Early Exits (EE)
Early Exits[57, 32, 29, 46], as a leading method in dynamic inference for computational saving, propose the termination of predictions as soon as the necessary computations are completed. While promising as a general model compression technique, research in this domain is still developing. Many implementations of EE are employed in dedicated designs to address efficiency[21, 25, 47, 66, 6, 19, 62]. Some other work explore the feasibility of introducing EE to general cases[33, 32, 29]. However, challenges remain in generalizing early exits, such as determining the optimal exit points and constructing efficient exits. Meanwhile, although not enabled by every work, the routing policy in early exits, is usually adjustable post-training and even during runtime, allowing users to finely tune the system’s behavior to meet specific requirements. DyCE[60] was introduced to search the best routing configurations, i.e. which exit to use, for different trade-off preferences in a given system. This allows for a range of configurations to suit varied demands. The research presented in this paper, however, tackles the latter challenge, focusing on exploring what are most efficient exits.
@@ 3.1.2 Efficiency
We can quantitatively evaluate the efficiency of TinySaver in an EE-based system with two exits. The dataset can be divided into two categories: (I) samples meeting the exit criteria and processed by the saver , and (II) those processed by the base model . Cat. I, filtered through exit criteria, is confidently predicted by , which generally indicates reliable performance. High confidence in predictions typically correlates with reliability, as evidenced in Fig. 3. Here, we examine varying confidence threshold values to chart the performance of both and base models for Cat. I samples. Initially, when is small the Cat. I samples are limited to very easy ones, both models exhibit very high performance on them. As the proportion of early exited samples increases, maintains a relatively high accuracy, while ’s performance declines more rapidly. Although large models generally outperform smaller models on average, this does not imply superiority in every individual sample. Given that these samples are selected based on ’s confidence, it’s not surprising that may perform better on Cat. I samples. ’s performance eventually diminishes as it handles more samples, but there remains a range where results for Cat. I samples are enhanced. Since Cat. II samples continue to be processed by , the system’s performance on this segment remains equivalent to that of the large model. Thus, improvements in Cat. I directly contribute to the overall enhancement of the system.
@@ 4 Experiment
| Base Model | Saver Model | Required FLOPs(%) when Acc. Drop | Max Performance | ||||||||||||||||||||
@@ 4.4 Comparison with related work
| Method | Model |
Box AP | GFLOPs | ||
| - | YOLOv8n[28] | IN1k | 37.3 | 9 | ||
| - | YOLOv8s[28] | IN1k | 44.9 | 29 | ||
| Efficient-DETR[69] | R50[24] | IN1k | 45.1 | 210 | ||
| TinySaver() |
IN1k | 45.3 | 71 | ||
| - | YOLOv8m[28] | IN1k | 50.2 | 79 | ||
| TinySaver() |
IN1k | 50.7 | 114 | ||
| Cascade Mask R-CNN[3] | Swin-T[40] | IN1k | 50.4 | 745 | ||
| TinySaver() |
Obj365[53], IN1k | 51.5 | 589 | ||
| Cascade Mask R-CNN[3] | Swin-S[40] | IN1k | 51.9 | 838 | ||
| - | YOLOv8l[28] | IN1k | 52.9 | 165 | ||
| Mask-RCNN[23] | ConvNeXt V2-B[64] | IN1k | 52.9 | 486 | ||
| TinySaver() |
Obj365[53], IN1k | 53.1 | 645 | ||
| RT-DETR[41] | R50[24] | IN1k | 53.1 | 136 | ||
| - | YOLOv8x[28] | IN1k | 53.9 | 258 | ||
| HTC++[5] | Swin-B[40] | IN22k | 56.4 | 1043 | ||
| TinySaver() |
Obj365[53], IN1k | 56.5 | 671 | ||
| HTC++[5] | Swin-L[40] | IN22k | 57.1 | 1470 | ||
| TinySaver() |
Obj365[53], IN1k | 60.3 | 1076 | ||
| Soft-teacher[68], HTC++[5] | Swin-L[40] | Obj365[53] | 60.1 | 1470 | ||
| TinySaver() |
Obj365[53], IN1k | 63.0 | 1497 | ||
| Co-DINO[74] | Swin-L[40] | Obj365[53] | 64.1 | 1527† |
@@ 5 Conclusion
Acknowledgments:
This publication has emanated from research conducted with the financial
support of Science Foundation Ireland under Grant number 18/CRT/6183.

### v2

@@ Abstract
This paper introduces TinySaver, an early-exit-like dynamic model compression approach which employs tiny models to substitute large models adaptively. Distinct from traditional compression techniques, dynamic methods like TinySaver can leverage the difficulty differences to allow certain inputs to complete their inference processes early, thereby conserving computational resources. Most existing early exit designs are implemented by attaching additional network branches to the model’s backbone. Our study, however, reveals that completely independent tiny models can replace a substantial portion of the larger models’ job with minimal impact on performance. Employing them as the first exit can remarkably enhance computational efficiency. By searching and employing the most appropriate tiny model as the computational saver for a given large model, the proposed approaches work as a novel and generic method to model compression. This finding will help the research community in exploring new compression methods to address the escalating computational demands posed by rapidly evolving AI models. Our evaluation of this approach in ImageNet-1k classification demonstrates its potential to reduce the number of compute operations by up to 90%, with only negligible losses in performance, across various modern vision models.
@@ 2.2.1 Early Exits (EE)
Early Exits[58, 33, 30, 47, 71], as a leading method in dynamic inference for computational saving, propose the termination of predictions as soon as the necessary computations are completed. While promising as a general model compression technique, research in this domain is still developing. Many implementations of EE are employed in dedicated designs to address efficiency[22, 26, 48, 68, 6, 19, 63, 21]. Some other work explore the feasibility of integrating EE with more cases[34, 33, 30, 64, 76]. However, challenges remain in generalizing early exits, such as determining the optimal exit points and constructing efficient exits. Meanwhile, although not enabled by every work, the routing policy in early exits, is usually adjustable post-training and even during runtime, allowing users to finely tune the system’s behavior to meet specific requirements. DyCE[61] was introduced to search the best routing configurations, i.e. which exit to use, for different trade-off preferences in real time. The research presented in this paper, however, tackles the latter challenge, focusing on exploring what are most efficient exits.
@@ 3.1.2 Efficiency
We can quantitatively evaluate the efficiency of TinySaver in an EE-based system with two exits. The dataset can be divided into two categories: (I) samples meeting the exit criteria and processed by the saver , and (II) those processed by the base model . Cat. I, filtered through exit criteria, is confidently predicted by , which generally indicates reliable performance. High confidence in predictions typically correlates with reliability, as evidenced in Fig. 3. Here, we examine varying confidence threshold values to chart the performance of both and base models for Cat. I samples. Initially, when is large the Cat. I samples are limited to very easy ones, both models exhibit very high performance on them. As the proportion of early exited samples increases, maintains a relatively high accuracy, while ’s performance declines more rapidly. Although large models generally outperform smaller models on average, this does not imply superiority in every individual sample. Given that these samples are selected based on ’s confidence, it’s not surprising that may perform better on Cat. I samples. ’s performance eventually diminishes as it handles more samples, but there remains a range where results for Cat. I samples are enhanced. Since Cat. II samples continue to be processed by , the system’s performance on this segment remains equivalent to that of the large model. Thus, improvements in Cat. I directly contribute to the overall enhancement of the system.
@@ 4 Experiment
4.1 Compression of general models
| Base Model | Saver Model | Required GFLOPs(%) when Acc. Drop | Max Performance | ||||||||||||||||||||
@@ 4.4 Comparison with related work
Method Model Pretrained Dataset Box AP GFLOPs - YOLOv8n[29] IN1k 37.3 9 - YOLOv8s[29] IN1k 44.9 29 Efficient-DETR[72] R50[25] IN1k 45.1 210 TinySaver() YOLOv8x[29] +YOLOv8n[29] IN1k 45.3 71 - YOLOv8m[29] IN1k 50.2 79 TinySaver() YOLOv8l[29] +YOLOv8s[29] IN1k 50.7 114 Cascade Mask R-CNN[3] Swin-T[41] IN1k 50.4 745 TinySaver() Co-DINO-Swin-L[78] + YOLOv8n[29] Obj365[54], IN1k 51.5 589 Cascade Mask R-CNN[3] Swin-S[41] IN1k 51.9 838 - YOLOv8l[29] IN1k 52.9 165 Mask-RCNN[24] ConvNeXt V2-B[66] IN1k 52.9 486 TinySaver() Co-DINO-Swin-L[78] +YOLOv8n[29] Obj365[54], IN1k 53.1 645 RT-DETR[42] R50[25] IN1k 53.1 136 - YOLOv8x[29] IN1k 53.9 258 HTC++[5] Swin-B[41] IN22k 56.4 1043 TinySaver() Co-DINO-Swin-L[78] +YOLOv8x[29] Obj365[54], IN1k 56.5 671 HTC++[5] Swin-L[41] IN22k 57.1 1470 TinySaver() Co-DINO-Swin-L[78] +YOLOv8n[29] Obj365[54], IN1k 60.3 1076 Soft-teacher[70], HTC++[5] Swin-L[41] Obj365[54] 60.1 1470 TinySaver() Co-DINO-Swin-L[78] +YOLOv8n[29] Obj365[54], IN1k 63.0 1497 Co-DINO[78] Swin-L[41] Obj365[54] 64.1 1527† †The FLOPs is measured by us, otherwise are collected from corresponding work.
@@ 5 Conclusion
Acknowledgment
This publication has emanated from research conducted with the financial support of Science Foundation Ireland under Grant number 18/CRT/6183.